Given P(A),find P (not A)

P(A)=3/8
Thank You for the straightfoward and correct answers you all send me.

Answer by nycsharkman(136)

Given P(A),find P (not A)
P(A)=3/8
I think P(not A) means to subtract 3/8 from 1.
So, 1 - 3/8 = 8/8 - 3/8 = 5/8
So, P(not A) = 5/8